# SEE: See Everything Every Time - Adaptive Brightness Adjustment for Broad Light Range Images via Events

### v1

@@ 4.2 Diversity Analysis
To analyze the diversity of our dataset, we evaluate the SEE-600K dataset in terms of both illumination and scene variety. Firstly, as shown in Fig.  4, for each scene, we capture multiple lighting conditions, including high-light, normal-light, and low-light, ensuring a broad range of illumination coverage. This approach enhances the dataset’s ability to support tasks requiring robust learning across diverse lighting variations. Secondly, the SEE-600K dataset includes 202 distinct scenes, as shown in Fig.  5, with a total of over 600K data samples, making it approximately 20 times larger than the latest SDE dataset [2]. This substantial increase in scale provides a richer foundation for training models with improved generalization capabilities. Additionally, we compare SEE-600K with several other event-based datasets, as shown in Tab.  1. Unlike previous datasets such as DVS-Dark [35] or EvLowLight [1], SEE-600K includes a broader range of illumination conditions. Furthermore, SEE-600K uniquely offers multiple groups of recordings per scene, enabling more robust learning of brightness transitions across various lighting conditions. This makes SEE-600K a significantly more comprehensive resource in terms of scale and diversity compared to other datasets like LIE [22], RLED [36], and SDE [2], which either lack diversity in lighting conditions or include fewer frames. Finally, we assess the diversity of the dataset using a word cloud generated by ChatGPT-4o [47], as shown in Fig. 5 (b). This word cloud illustrates the variety of scene categories and objects present in the dataset, further emphasizing its comprehensive coverage. With scenes ranging from urban environments to industrial settings and natural landscapes, SEE-600K captures a diverse array of objects and contexts, making it a valuable resource for event-based imaging research.
@@ 5 Methods
Overview: As shown in Fig. 6, our framework, SEE-Net, consists of four implementation parts: (a) Inputs and Preprocessing, (b) Encoding, (c) Decoding, and (d) Loss Function. The input is an image $I_{i}$ and its corresponding events $E$. The output is a brightness-adjustable image $I_{o}$, where the brightness is controlled by the prompt $B\in(0,1)$. During training, the brightness prompt $B$ is calculated according to the target image. On the other hand, during testing, $B$ can be freely set, with a default value of 0.5, which follows the exposure control constraint [39, 48]. Overall, the SEE-Net $f_{see}$ can be described by the Eq. 2 to match our learning task in Sec. 3.
@@
Inputs and Processing: This part aims to transform initial inputs into features that retain original information for the encoding stage. The inputs consist of the image $I_{i}$ and the events $E$, where $I_{i}$ has a dimension of $H\times W\times 3$ (with $H$ and $W$ representing the height and width, and 3 representing the color channel number). The event stream $E$ is represented as a voxel grid [49] with a dimension of $H\times W\times M$, where $M$ represents the number of time slices of events. The events include color information [50], which was overlooked in previous works, e.g., [2, 1]. Specifically, this DVS346 sensor records events with Bayer Pattern [51]. To effectively embed both the color and positional information during framework training, we design the position and Bayer Pattern embeddings, as shown in Fig. 6 (a). The position and Bayer Pattern are denoted as a vector $(x,y,bp)$, where $x,y$ is the pixel position, and $bp$ denotes the Bayer Pattern index, which takes a value from 0 to 3. We embed this vector into a higher-dimensional feature, termed as $P$, and concatenate it with the inputs. Two layers $1\times 1$ convolutions, denote $f_{e}$ and $f_{i}$, are then applied to obtain the initial event features $F_{e}$ and image features $F_{i}$. This process is described by the Eq. 3, where $f_{cat}$ denotes the concatenation function.
@@
[…]
@@
DCE [52] 0.66 0.01 13.91 0.2659 13.38 0.1842 13.64 0.2250
SNR [53] 26.35 4.01 20.05 0.6302 22.18 0.6611 21.12 0.6457
UFormer [54] 12.00 5.29 21.09 0.7524 22.32 0.7469 21.71 0.7497
LLFlow [55] 409.50 39.91 20.92 0.6610 21.68 0.6467 21.30 0.6539
Retinexformer [23] 15.57 1.61 21.30 0.6920 22.92 0.6834 22.11 0.6877
E2VID+ [13] 27.99 10.71 15.19 0.5891 15.01 0.5765 15.10 0.5828
ELIE [22] 440.32 33.36 19.98 0.6168 20.69 0.6533 20.34 0.6350
[…]
[56] 170.32 7.38 21.29 0.6786 22.08 0.7052 21.68 0.6919
eSL-Net [57] 560.94 0.56 21.25 0.7277 22.42 0.7187 21.84 0.7232
[21] 44.71 47.06 21.79 0.7051 23.35 0.6895 22.57 0.6973
EvLowlight [1] 524.95 15.49 20.57 0.6217 20.04 0.6485 20.31 0.6351
EvLight [2] 180.90 22.73 22.44 0.7697 23.21 0.7505 22.83 0.7601
@@
Encoding: In this stage, we aim to obtain the BLR by employing the event feature $F_{e}$ to enhance the image feature $F_{i}$, facilitating noise reduction and the acquisition of broader light range information. Since $F_{e}$ contains rich information about the lighting changes across different intensity levels, we use it as the source for representing the broader light range. However, event data only records changes in illumination, which differ fundamentally from the static RGB frame modality. This makes directly utilizing event data for broader light representation challenging. To address this, we employ a cross-attention [58] for feature fusion, producing the initial fused broad-spectrum feature $F_{1}$, expressed as $F_{1}=f_{c_{0}}(F_{e},F_{i})$, where $f_{c_{0}}$ is a cross-attention block. Then, inspired by previous works [57], we utilize sparse learning to generate residuals for $F_{1}$ from the event features $F_{e}$. These residuals are progressively generated from the loop that executes $L$ times. Multiple iterations are used because they allow the model to iteratively refine the residuals, capturing finer details and enhancing the feature representations by progressively integrating information from the events. A single loop of this process can be expressed as, $F_{j+1}=f_{l}(F_{e},F_{1},F_{j})$, where $f_{l}$ is a loop function that contains two cross-attention blocks as shown in Fig. 6 (b), where $F_{j}$ and $F_{j+1}$ are the input and output of one loop. After $L$ iterations, the final feature $F_{L}$ represents the BLR, as described by Eq. 4.
@@
Decoding: The objective of this part is to decode the BLR into a brightness-adjustable image $I_{o}$. In designing this decoder, we focus on two key insights: (1) The decoding process should be pixel-wise and efficient, allowing for greater flexibility during model deployment; (2) The embedding of the brightness information should be thorough and fully integrated. With these insights, we design the decoder with only a 5-layer MLP as shown in Fig 6 (c). Our decoder begins by encoding the brightness prompt $B\in(0,1)$ into an embedding vector. To effectively encode the high-frequency brightness prompt into features that are easier for the network to learn [59], we introduce a learnable embedding, denoted as $\bm{B}=f_{pe}(B)=f_{mlp}(f_{cat}(f_{mlp}(B),B))$, which consists of two MLP layers. Through this embedding, the brightness prompt $B$ is transformed into a vector $\bm{B}$, matching the dimensions of the BLR channels. We then integrate this embedding $\bm{B}$ into the decoder. To ensure the brightness prompt is fully incorporated and prevent information loss through multiple MLP layers, we employ a multi-step embedding approach, as displayed in Eq. 5, which guarantees that the brightness is progressively embedded throughout the decoding process. During the training phase, the prompt $B$ is derived from the reference image by applying $f_{b}$ to calculate the global average brightness. In contrast, during the testing phase, $B$ can be set freely, with a typical example being a value of 0.5.
@@
Loss Function: The purpose of our loss function is to supervise the prediction $I_{o}$ using the ground truth $I_{t}$, with the corresponding brightness $B=f_{b}(I_{t})$. The loss function consists of two main components: image reconstruction loss $\mathcal{L}_{i}$ and gradient loss $\mathcal{L}_{g}$. The image reconstruction loss is Charbonnier loss [60], which effectively handles both small and large errors. Additionally, we employ gradient loss to improve the structural consistency of the output image. This is achieved by enforcing $L_{1}$ constraints on the gradients of both the output and ground truth images. Therefore, the overall loss function is formulated as a weighted sum of the image loss and gradient loss, as exhibited in Eq. 6. Here, $\nabla$ denotes the gradient operator, and $\lambda_{1}$ and $\lambda_{2}$ are the weights that balance the contributions of two loss terms.
@@ 6 Experiments
Training Dataset Methods low light high light normal light
PSNR SSIM $L_{1}$ PSNR SSIM $L_{1}$ PSNR SSIM $L_{1}$
SDE DCE [52] 9.10 0.0968 0.3572 6.26 0.3419 0.4649 10.79 0.3992 0.2524
eSL Net [57] 11.92 0.3275 0.2703 6.66 0.1672 0.4001 7.65 0.2685 0.3481
[21] 12.41 0.4001 0.2487 5.53 0.1950 0.4534 6.58 0.2805 0.4129
EvLowLight [1] 12.68 0.4341 0.2338 4.11 0.3071 0.6062 7.01 0.3950 0.4520
EvLight [2] 13.07 0.4651 0.2337 5.12 0.1005 0.4842 6.29 0.2805 0.4336
@@
SEE eSL Net [57] 11.95 0.3845 0.2421 12.84 0.4660 0.2076 13.45 0.5682 0.1957
EvLowLight [1] 12.83 0.4511 0.2151 12.79 0.4696 0.2084 13.04 0.5531 0.2144
[21] 13.48 0.5068 0.1946 12.30 0.4766 0.2221 13.70 0.5474 0.2151
EvLight [2] 13.70 0.5150 0.1960 13.45 0.4918 0.1990 13.63 0.5924 0.2004
@@
Implementation Details: Our experiments use the Adam optimizer with an initial learning rate of $42e-$ 4 for all the experiments. We train our model for 40 epochs on the SDE dataset [2]. On the SEE-600K dataset, we train for only 20 epochs, as SEE-600K is extremely large. All of our training is conducted on an HPC cluster, with a batch size of 2. To enhance data diversity, we apply random cropping to the images and perform random flips and rotations.
Evaluation Metrics: We maintain consistency with previous methods [2, 1] by using PSNR and SSIM [61]. However, since our proposed new problem is highly challenging and most current approaches perform poorly on our SEE-600K dataset, we additionally introduce the $L_{1}$ distance as a reference.
Dataset: We conduct experiments on two real-world datasets: (1) SDE [2] comprises 91 scenes, with 76 for training and 15 for testing. Each scene includes a pair of low-light and normal-light images along with their corresponding events. (2) SEE-600K consists of 202 scenes, with each scene containing an average of four sets of videos under different lighting conditions, ranging from low light to bright light. During each training session, we randomly select one set of normal-light images as the reference and use the remaining sets as inputs. For example, for one scene with one low-light, two normal-light, and one high-light set, we generate six pairs of training data.
@@ 6.1 Comparisons Experiments
Comparative Methods: We categorize the competitive approaches into four groups. Firstly, DCE [52] is a classical approach that can adjust the image brightness curve to achieve normal lighting. Secondly, there are strategies that only use images as input, including SNR [53], UFormer [54], LLFlow [55], and RetinexFormer [23]. Thirdly, we consider methods that rely solely on events, e.g., E2VID+ [13]. Tertiary, we examine event-guided low-light enhancement frameworks. This group includes single-frame input methods, e.g., eSL-Net [57], [21], [56]and EvLight [2], as well as multi-frame input strategies like EvLowLight [1]. Furthermore, we also compared the HDR reconstruction method HDRev [16]. We retrain all methods, following the open-source code when available; for approaches without open-source code, we replicate them based on their respective papers.
@@
(1) performance limitations of single-modal methods: Methods utilizing only a single modality exhibit limited performance, as shown in Tab. 2. This trend highlights the need to integrate both modalities to achieve improved results, as demonstrated by methods like DCE [52], SRN [53], and UFormer [54] in Tab.2. The employment of a single modality does not supplement the dynamic range information, leading to relatively poorer performance, as shown in Fig.7.
@@
(3) impact of indoor and outdoor conditions: Notably, performance in low-light indoor scenarios is inferior to that in outdoor settings, as shown in Fig. 7 (I) and (III). This discrepancy may be attributed to the issues of flickering light sources commonly found indoors [62]. Our SEE-Net consistently achieves the best results across both scenarios, with a model size of just $1.9M$$1/10$ parameter count of other SOTA methods—demonstrating its efficiency and compactness in low-light image enhancement.
@@
(1) Models trained on SDE: Models trained on the SDE dataset perform reasonably well when tested on the SEE-600K dataset, particularly under low-light conditions. Notably, the DCE [52] method achieves the best performance in high-light scenarios, demonstrating its strong generalization capability, particularly for self-supervised learning approaches. However, in high-light and normal-light conditions, almost all methods fail, which highlights a significant limitation. This indicates that models trained on the SDE dataset learn a fixed multiplicative enhancement for illumination, and are not adaptable to a wider range of lighting conditions. This emphasizes the importance of redefining low-light enhancement as an illumination adjustment problem, as we have proposed.

### v2

@@ 4.2 Diversity Analysis
To analyze the diversity of our dataset, we evaluate the SEE-600K dataset in terms of both illumination and scene variety. Firstly, as shown in Fig.  4, for each scene, we capture multiple lighting conditions, including high-light, normal-light, and low-light, ensuring a broad range of illumination coverage. This approach enhances the dataset’s ability to support tasks requiring robust learning across diverse lighting variations. Secondly, the SEE-600K dataset includes 202 distinct scenes, as shown in Fig.  5, with a total of over 600K data samples, making it approximately 20 times larger than the latest SDE dataset (Liang et al, 2024). This substantial increase in scale provides a richer foundation for training models with improved generalization capabilities. Additionally, we compare SEE-600K with several other event-based datasets, as shown in Tab.  1. Unlike previous datasets such as DVS-Dark (Zhang et al, 2020) or EvLowLight (Liang et al, 2023), SEE-600K includes a broader range of illumination conditions. Furthermore, SEE-600K uniquely offers multiple groups of recordings per scene, enabling more robust learning of brightness transitions across various lighting conditions. This makes SEE-600K a significantly more comprehensive resource in terms of scale and diversity compared to other datasets like LIE (Jiang et al, 2023), RLED (Liu et al, 2024), and SDE (Liang et al, 2024), which either lack diversity in lighting conditions or include fewer frames. Finally, we assess the diversity of the dataset using a word cloud generated by ChatGPT-4o (Hernández-Flores et al, 2025), as shown in Fig. 5 (b). This word cloud illustrates the variety of scene categories and objects present in the dataset, further emphasizing its comprehensive coverage. With scenes ranging from urban environments to industrial settings and natural landscapes, SEE-600K captures a diverse array of objects and contexts, making it a valuable resource for event-based imaging research.
@@ 5 Methods
Overview: As shown in Fig. 6, our framework, SEE-Net, consists of four implementation parts: (a) Inputs and Preprocessing, (b) Encoding, (c) Decoding, and (d) Loss Function. The input is an image $I_{i}$ and its corresponding events $E$. The output is a brightness-adjustable image $I_{o}$, where the brightness is controlled by the prompt $B\in(0,1)$. During training, the brightness prompt $B$ is calculated according to the target image. On the other hand, during testing, $B$ can be freely set, with a default value of 0.5, which follows the exposure control constraint (Mertens et al, 2009, 2007). Overall, the SEE-Net $f_{see}$ can be described by the Eq. 2 to match our learning task in Sec. 3.
@@
Inputs and Processing: This part aims to transform initial inputs into features that retain original information for the encoding stage. The inputs consist of the image $I_{i}$ and the events $E$, where $I_{i}$ has a dimension of $H\times W\times 3$ (with $H$ and $W$ representing the height and width, and 3 representing the color channel number). The event stream $E$ is represented as a voxel grid (Tulyakov et al, 2022) with a dimension of $H\times W\times M$, where $M$ represents the number of time slices of events. The events include color information Scheerlinck et al (2019), which was overlooked in previous works, e.g., (Liang et al, 2024, 2023). Specifically, this DVS346 sensor records events with Bayer Pattern (Lukac et al, 2005). To effectively embed both the color and positional information during framework training, we design the position and Bayer Pattern embeddings, as shown in Fig. 6 (a). The position and Bayer Pattern are denoted as a vector $(x,y,bp)$, where $x,y$ is the pixel position, and $bp$ denotes the Bayer Pattern index, which takes a value from 0 to 3. We embed this vector into a higher-dimensional feature, termed as $P$, and concatenate it with the inputs. Two layers $1\times 1$ convolutions, denote $f_{e}$ and $f_{i}$, are then applied to obtain the initial event features $F_{e}$ and image features $F_{i}$. This process is described by the Eq. 3, where $f_{cat}$ denotes the concatenation function.
@@
Table 2: Comparison of different methods on the SDE dataset. The best performances is highlighted in bold. $\dagger$} refers to the original model for the HDR task, which is fine-tuned and trained on SDE
@@
DCE (Guo et al, 2020) 0.66 0.01 13.91 0.2659 13.38 0.1842 13.64 0.2250
SNR (Xu et al, 2022) 26.35 4.01 20.05 0.6302 22.18 0.6611 21.12 0.6457
UFormer (Wang et al, 2022) 12.00 5.29 21.09 0.7524 22.32 0.7469 21.71 0.7497
LLFlow (Wu et al, 2023) 409.50 39.91 20.92 0.6610 21.68 0.6467 21.30 0.6539
Retinexformer (Cai et al, 2023) 15.57 1.61 21.30 0.6920 22.92 0.6834 22.11 0.6877
E2VID+ (Stoffregen et al, 2020) 27.99 10.71 15.19 0.5891 15.01 0.5765 15.10 0.5828
ELIE (Jiang et al, 2023) 440.32 33.36 19.98 0.6168 20.69 0.6533 20.34 0.6350
HDRev(Yang et al, 2023) $\dagger$} 118.65 13.42 21.13 0.6239 21.82 0.6824 21.47 0.6531
(Wang et al, 2023a) 170.32 7.38 21.29 0.6786 22.08 0.7052 21.68 0.6919
eSL-Net (Wang et al, 2020) 560.94 0.56 21.25 0.7277 22.42 0.7187 21.84 0.7232
(Liu et al, 2023) 44.71 47.06 21.79 0.7051 23.35 0.6895 22.57 0.6973
EvLowlight (Liang et al, 2023) 524.95 15.49 20.57 0.6217 20.04 0.6485 20.31 0.6351
EvLight (Liang et al, 2024) 180.90 22.73 22.44 0.7697 23.21 0.7505 22.83 0.7601
@@
Encoding: In this stage, we aim to obtain the BLR by employing the event feature $F_{e}$ to enhance the image feature $F_{i}$, facilitating noise reduction and the acquisition of broader light range information. Since $F_{e}$ contains rich information about the lighting changes across different intensity levels, we use it as the source for representing the broader light range. However, event data only records changes in illumination, which differ fundamentally from the static RGB frame modality. This makes directly utilizing event data for broader light representation challenging. To address this, we employ a cross-attention (Liang et al, 2021) for feature fusion, producing the initial fused broad-spectrum feature $F_{1}$, expressed as $F_{1}=f_{c_{0}}(F_{e},F_{i})$, where $f_{c_{0}}$ is a cross-attention block. Then, inspired by previous works (Wang et al, 2020), we utilize sparse learning to generate residuals for $F_{1}$ from the event features $F_{e}$. These residuals are progressively generated from the loop that executes $L$ times. Multiple iterations are used because they allow the model to iteratively refine the residuals, capturing finer details and enhancing the feature representations by progressively integrating information from the events. A single loop of this process can be expressed as, $F_{j+1}=f_{l}(F_{e},F_{1},F_{j})$, where $f_{l}$ is a loop function that contains two cross-attention blocks as shown in Fig. 6 (b), where $F_{j}$ and $F_{j+1}$ are the input and output of one loop. After $L$ iterations, the final feature $F_{L}$ represents the BLR, as described by Eq. 4.
@@
Decoding: The objective of this part is to decode the BLR into a brightness-adjustable image $I_{o}$. In designing this decoder, we focus on two key insights: (1) The decoding process should be pixel-wise and efficient, allowing for greater flexibility during model deployment; (2) The embedding of the brightness information should be thorough and fully integrated. With these insights, we design the decoder with only a 5-layer MLP as shown in Fig 6 (c). Our decoder begins by encoding the brightness prompt $B\in(0,1)$ into an embedding vector. To effectively encode the high-frequency brightness prompt into features that are easier for the network to learn (Vaswani, 2017), we introduce a learnable embedding, denoted as $\bm{B}=f_{pe}(B)=f_{mlp}(f_{cat}(f_{mlp}(B),B))$, which consists of two MLP layers. Through this embedding, the brightness prompt $B$ is transformed into a vector $\bm{B}$, matching the dimensions of the BLR channels. We then integrate this embedding $\bm{B}$ into the decoder. To ensure the brightness prompt is fully incorporated and prevent information loss through multiple MLP layers, we employ a multi-step embedding approach, as displayed in Eq. 5, which guarantees that the brightness is progressively embedded throughout the decoding process. During the training phase, the prompt $B$ is derived from the reference image by applying $f_{b}$ to calculate the global average brightness. In contrast, during the testing phase, $B$ can be set freely, with a typical example being a value of 0.5.
@@
Loss Function: The purpose of our loss function is to supervise the prediction $I_{o}$ using the ground truth $I_{t}$, with the corresponding brightness $B=f_{b}(I_{t})$. The loss function consists of two main components: image reconstruction loss $\mathcal{L}_{i}$ and gradient loss $\mathcal{L}_{g}$. The image reconstruction loss is Charbonnier loss (Lai et al, 2018), which effectively handles both small and large errors. Additionally, we employ gradient loss to improve the structural consistency of the output image. This is achieved by enforcing $L_{1}$ constraints on the gradients of both the output and ground truth images. Therefore, the overall loss function is formulated as a weighted sum of the image loss and gradient loss, as exhibited in Eq. 6. Here, $\nabla$ denotes the gradient operator, and $\lambda_{1}$ and $\lambda_{2}$ are the weights that balance the contributions of two loss terms.
@@ 6 Experiments
Training
Dataset
Methods low light high light normal light
PSNR
SSIM
$L_{1}$
PSNR
SSIM
$L_{1}$
PSNR
SSIM
$L_{1}$
SDE DCE (Guo et al, 2020) 9.10 0.0968 0.3572 6.26 0.3419 0.4649 10.79 0.3992 0.2524
eSL Net (Wang et al, 2020) 11.92 0.3275 0.2703 6.66 0.1672 0.4001 7.65 0.2685 0.3481
Liu et al (2023) 12.41 0.4001 0.2487 5.53 0.1950 0.4534 6.58 0.2805 0.4129
EvLowLight (Liang et al, 2023) 12.68 0.4341 0.2338 4.11 0.3071 0.6062 7.01 0.3950 0.4520
EvLight (Liang et al, 2024) 13.07 0.4651 0.2337 5.12 0.1005 0.4842 6.29 0.2805 0.4336
@@
SEE eSL Net (Wang et al, 2020) 11.95 0.3845 0.2421 12.84 0.4660 0.2076 13.45 0.5682 0.1957
EvLowLight (Liang et al, 2023) 12.83 0.4511 0.2151 12.79 0.4696 0.2084 13.04 0.5531 0.2144
Liu et al (2023) 13.48 0.5068 0.1946 12.30 0.4766 0.2221 13.70 0.5474 0.2151
EvLight (Liang et al, 2024) 13.70 0.5150 0.1960 13.45 0.4918 0.1990 13.63 0.5924 0.2004
@@
Implementation Details: Our experiments use the Adam optimizer with an initial learning rate of 2e42e-4 for all the experiments. We train our model for 40 epochs on the SDE dataset (Liang et al, 2024). On the SEE-600K dataset, we train for only 20 epochs, as SEE-600K is extremely large. All of our training is conducted on an HPC cluster, with a batch size of 2. To enhance data diversity, we apply random cropping to the images and perform random flips and rotations.
Evaluation Metrics: We maintain consistency with previous methods (Liang et al, 2024, 2023) by using PSNR and SSIM (Wang et al, 2004). However, since our proposed new problem is highly challenging and most current approaches perform poorly on our SEE-600K dataset, we additionally introduce the $L_{1}$ distance as a reference.
Dataset: We conduct experiments on two real-world datasets: (1) SDE (Liang et al, 2024) comprises 91 scenes, with 76 for training and 15 for testing. Each scene includes a pair of low-light and normal-light images along with their corresponding events. (2) SEE-600K consists of 202 scenes, with each scene containing an average of four sets of videos under different lighting conditions, ranging from low light to bright light. During each training session, we randomly select one set of normal-light images as the reference and use the remaining sets as inputs. For example, for one scene with one low-light, two normal-light, and one high-light set, we generate six pairs of training data.
@@ 6.1 Comparisons Experiments
Comparative Methods: We categorize the competitive approaches into four groups. Firstly, DCE (Guo et al, 2020) is a classical approach that can adjust the image brightness curve to achieve normal lighting. Secondly, there are strategies that only use images as input, including SNR (Xu et al, 2022), UFormer (Wang et al, 2022), LLFlow (Wu et al, 2023), and RetinexFormer (Cai et al, 2023). Thirdly, we consider methods that rely solely on events, e.g., E2VID+ (Stoffregen et al, 2020). Tertiary, we examine event-guided low-light enhancement frameworks. This group includes single-frame input methods, e.g., eSL-Net (Wang et al, 2020), (Liu et al, 2023), (Wang et al, 2023a)and EvLight (Liang et al, 2024), as well as multi-frame input strategies like EvLowLight (Liang et al, 2023). Furthermore, we also compared the HDR reconstruction method HDRev (Yang et al, 2023). We retrain all methods, following the open-source code when available; for approaches without open-source code, we replicate them based on their respective papers.
@@
(1) performance limitations of single-modal methods: Methods utilizing only a single modality exhibit limited performance, as shown in Tab. 2. This trend highlights the need to integrate both modalities to achieve improved results, as demonstrated by methods like DCE (Guo et al, 2020), SRN (Xu et al, 2022), and UFormer (Wang et al, 2022) in Tab.2. The employment of a single modality does not supplement the dynamic range information, leading to relatively poorer performance, as shown in Fig.7.
@@
(3) impact of indoor and outdoor conditions: Notably, performance in low-light indoor scenarios is inferior to that in outdoor settings, as shown in Fig. 7 (I) and (III). This discrepancy may be attributed to the issues of flickering light sources commonly found indoors (Xu et al, 2023a). Our SEE-Net consistently achieves the best results across both scenarios, with a model size of just 1.9M1.9M—10% parameter count of other SOTA methods—demonstrating its efficiency and compactness in low-light image enhancement.
@@
(1) Models trained on SDE: Models trained on the SDE dataset perform reasonably well when tested on the SEE-600K dataset, particularly under low-light conditions. Notably, the DCE (Guo et al, 2020) method achieves the best performance in high-light scenarios, demonstrating its strong generalization capability, particularly for self-supervised learning approaches. However, in high-light and normal-light conditions, almost all methods fail, which highlights a significant limitation. This indicates that models trained on the SDE dataset learn a fixed multiplicative enhancement for illumination, and are not adaptable to a wider range of lighting conditions. This emphasizes the importance of redefining low-light enhancement as an illumination adjustment problem, as we have proposed.